OpenStudy (anonymous):

Joyce paid $91.00 for an item at the store that was 30 percent off the original price. What was the original price?

OpenStudy (tkhunny):

Name the original price, please. Seriously, just give it a name.

OpenStudy (anonymous):

this is the exact question from my homework...the original price isn't given.

OpenStudy (tkhunny):

You didn't do what I asked. I did NOT say "what is the price", I said simply "name it". It needs a name. Give it one. Then we can talk about it. Seriously, just give it a name.

OpenStudy (anonymous):

what needs a name?? The item? I don't understand what you're asking.

OpenStudy (babalooo):

lol. okay. Joyce paid $91.00 for an item at the store that was 30 percent off the original price. What was the original price? so we know original x 30/100 = discount so original= discount+91(amount paid) which in other words, original x 70/100 = 91 (amount paid get it?

OpenStudy (babalooo):

@tkbunny is just asking you to give original price a unknown variable I believe like x or somethign

OpenStudy (anonymous):

@babalooo Got it! Thank you for the help!

OpenStudy (tkhunny):

There, see! @babalooo NAMED the original price "original". See, I told you that wasn't hard. It also nearly solves the problem. What do you not understand about "Name the original price, please." It's not tricky. Just name it.

OpenStudy (babalooo):

Please press the best response button to give me a metal if you think I was helpful. thanks.

OpenStudy (triciaal):

another approach 30% off means paid only 70% remember percent means per 100 70/100*original price= 91 solve for original price = 91*100/70